# FG-PE: Factor-graph Approach for Multi-robot Pursuit-Evasion

### v1

@@ III-C1 Dynamic Factor of Evader
$e_{t}^{qd}(x_{t}^{q})$ $={||x_{t}^{q}-x_{t-1}^{q}||}_{\Omega_{t}^{q}}^{2}~{},$ (5)
In Eq. (5), $x^{q}_{t}$ represents the evader’s position at time $t$. The cost function measures the prediction error of the relative transformation between two consecutive evader poses, comparing the predicted motion based on current pose estimates with the actual odometry measurements. This error is weighted by the uncertainty of those measurements.
III-C2 Dynamic Factor of Pursuers
@@
III-C3 Measurement factor between each pursuer and evader
@@
III-C4 Measurement factor between pursuers and obstacles
Pursuers need to localize themselves in the scene, and as a result, they will take measurements with respect to obstacles. In this factor, the pose of each pursuer is compared against the estimated pursuer’s pose at each time step. The measurement factor compares the estimated pose of the evader and the measured value. This measurement factor can be defined as follows:
$e_{t}^{m_{j}o_{i}}(x^{o},x^{p}_{t})$ $={||\hat{x}^{p}_{t}-h({x}^{o},{x}^{p}_{t})||}_{\Omega_{t}^{qo}}^{2}~{},$ (8)
where $\hat{x}^{p}_{t}$ represents the estimated position of a pursuer at time $t$, and $x^{o}$ represents the position of obstacle $o$.
III-C5 Planning factor for pursuers
@@
$e_{t}^{p_{j}mov}(\hat{x}^{q}_{t},\hat{x}_{t}^{p})$ $={||\hat{x}^{q}_{t}-\hat{x}_{t}^{p}||}_{\Omega_{t}^{pq}}^{2}~{}.$ (9)
III-C6 Collision Avoidance factor for pursuers
@@
where the objective is to define a safety bubble for each pursuer. The radius of the safety bubble is given by $d_{s}$. If the distance between pursuers $p_{i}$ and $p_{j}$ is less than $d_{s}$, the error term $||1-\frac{d_{p}}{c_{1}}||$ is applied. Here, $c_{1}$ is a constant, and $d_{p}$ represents the distance between the two pursuers $p_{i}$ and $p_{j}$.
III-C7 Obstacle Avoidance factor for pursuers
@@
Here, $d_{o}$ represents the distance between each moving pursuer and the obstacles in the scene, while $c_{2}$ is a constant.
The optimized cost function generates the path based on the estimated position of pursuers ($\hat{x}^{p}_{t}$) and the evader ($\hat{x}^{q}_{t}$) at time $t$.
III-D Time Complexity Analysis
To analyze the complexity of the problem, the worst-case scenario is considered. The number of variables and factors grows over time. The number of variables added at each time step is $N_{p}+N_{q}$. To investigate the number of factors added at each time step, the effect of each factor must be investigated. The number of factors added at each time step for the dynamic factor of pursuers $f^{dp}$, planning factor of pursuers $f^{pmove}$, and measurement factor between pursuers and evader $f^{mp}$ is the same and is $N_{p}$. The number of factors added for the dynamic factor of evader ($f^{dq}$) at each time step is $N_{q}$. The number of factors added at each time step for collision avoidance factor is $\frac{N_{p}\times(N_{p}-1)}{2}$. The number of factors added for the measurement factor between pursuers and obstacles ($f^{mo}$) and obstacle avoidance factor$(f^{op})$ is $N_{p}\times N_{o}$. The overall complexity is
@@ III-D Time Complexity Analysis
Figure 2: The number of different factors at each time step is shown in the figure above. In this figure, $f^{op}$ and $f^{mo}$ are increasing at the same rate, while $f^{dp}$, $f^{mp}$, and $f^{pmove}$ are increasing at the same speed.
As shown in Fig. 2, the number of measurement factors between pursuers and obstacles and obstacle avoidance factors increase faster compared with other factors and variables.
The LM solver can be approximated by $\widetilde{O}(\epsilon^{-2})$ [32] where $\widetilde{O}$ exhibits the existence of logarithmic factors in $\epsilon$, and $\epsilon$ represents an $\epsilon$-stationary point. By considering $m$ as solver parameters and considering the post-optimization that takes place in $O(I)$, the overall complexity would be $\widetilde{O}(n\epsilon^{-2}(n^{2}+m)+nI)$.
III-E Convergence Analysis
@@
$f_{i}(X_{i})\propto\exp\left\{-\frac{1}{2}\left\|h_{i}(X_{i})-z_{i}\right\|_{\Sigma_{i}}^{2}\right\}$ (13)
@@
$X^{\text{MAP}}=\arg\min_{X}\sum_{i}\left\|h_{i}(X_{i})-z_{i}\right\|_{\Omega_{i}}^{2}$ (14)
The LM optimizer can solve Eq. (14) by liniarizing it as:
$\Delta^{*}=\arg\min_{\Delta}||A\Delta-b||_{2}^{2}$ (15)
In Eq. (15), $A$ is the single measurement Jacobian, and $b$ shows all the prediction errors in the form of $z_{i}-h_{i}(X^{0}_{t})$. As a simple scenario, if there are two pursuers, $p_{1}$ and $p_{2}$, and they observe the evader only once the factor graph simplifies as shown in Fig. 3. This graph has a tree structure, and it converges surely [34]. In the factor graph, the factor function for the movements of pursuers is denoted as $f_{1}^{dp_{1}}$ for $p_{1}$ at time-step 0, as shown in Fig. 3. The prior factor of the pursuer $p$ is represented as $f_{0}^{p_{1}}$ at time-step 0.
Variable elimination is an exact inference method.
As an example, the probability function of variable $x_{3}^{q}$ can be found by summing out all other variables in the joint distribution:
@@
$\sum_{x_{0}^{p_{2}}}\cdots\sum_{x_{3}^{p_{2}}}P(x^{q}_{0},\ldots,x^{q}_{3},x_{0}^{p_{1}},\ldots,x_{3}^{p_{1}},x_{0}^{p_{2}},\ldots,x_{3}^{p_{2}})$ (16)
In Eq. (16), all variables are marginalized out, except $x_{3}^{q}$. By starting the elimination process in the graph shown in Fig. 3, we begin by eliminating node $x_{0}^{p_{1}}$. The involved factors are $f_{0}^{p_{1}}$, ${f_{0}^{mp_{1}}}$, and ${f_{1}^{dp_{1}}}$. After eliminating $x_{0}^{p_{1}}$, two marginal factors are created.
@@
III-F Uncertainty minimization analysis
@@
$x_{t}^{q}=x_{t-1}^{q}-(J^{T}J+\lambda I)^{-1}J^{T}e_{t}^{qd}$ (24)
In the above equation, $J$ represents the Jacobian matrix of the cost function, and $\lambda$ is the damping factor that balances between the gradient descent and Gauss-Newton methods.
By considering that the only factor existing in the problem is the dynamic factor, the uncertainty can be expressed as:
$\sigma_{T}=\sigma_{0}+\sum_{i=0}^{T}e^{q}_{t}-\Delta_{LM}$ (25)
@@
IV Results
Experiments are conducted in two environments: (i) a simulated environment and (ii) a real-world hardware environment. The performance analysis is based on the average distance traveled by pursuers, time, and the number of time steps. Parameters are set based on the robot web paper by Muraiet al. [26].
  • Dynamic factor: In this problem, the uncertainties for each step along the x-axis and y-axis are denoted as $\sigma_{x}=0.1$ and $\sigma_{y}=0.1$, respectively. Additionally, the uncertainty in orientation is $\sigma_{\theta}=0.01$.
@@ IV Results
  • Collision avoidance factor: In order to prevent pursuers from having collisions, this factor is utilized. The uncertainties for each step along the x-axis and y-axis are denoted as $\sigma_{x}$ = 1 and $\sigma_{y}$ = 0.1, respectively.
@@
  • Obstacle avoidance factor: In order to prevent pursuers having collision with obstacles, this factor is utilized. The uncertainties for each step along the x-axis and y-axis are denoted as $\sigma_{x}$ = 1 and $\sigma_{y}$= 0.1 respectively.
IV-A Simulation
@@
The simulated environment is developed in a way that the number of pursuers, evaders, and obstacles in the scene can be modified. The simulated environment is developed in a $3500\times 3500$ window size. Additionally, the simulated environment is capable of simulating different trajectories of the evader. The simulation is implemented in C++ and OpenCV library is utilised for visualization purpose.
IV-A1 Comparison
Various approaches have been explored in pursuit of strategies for capturing and chasing evaders. Among these, Qin et al. [35] introduced a hierarchical network-based model aimed at devising a cooperative pursuit strategy.
@@ IV-A1 Comparison
To ensure a fair comparison, the initial poses and speeds of both the evader and pursuers are kept the same for the proposed and compared methods. Two different variations of [35] are considered: Constant Bearing (CB) and Pure Pursuit (PP) strategies. In the CB strategy, the bearing angle between a pursuer and the evader remains fixed until the time of capture. Conversely, in the PP strategy, the pursuer’s velocity vector aligns with the line of sight. Our proposed method operates at a frequency of 20 Hz on average, while the compared methods operate at a frequency of 30 Hz on average.
@@
Figure 4: Average distance traveled by pursuers (top) and time to capture the evader (down) for different approaches. $V_{p}$ and $V_{q}$ show the rate of pursuer speed in comparison to evader speed. PP and CB methods are not able to catch the evader at all speed rates, in which no values are reported for them.
As can be seen in Fig. 4, the proposed method captures the evader faster in all speed rates that were tested, and also the average distance traveled by pursuers is much less in our method. In speed rates ($V_{p}$=1.05:$V_{q}$=1) and ($V_{p}=0.9:V_{q}=1$) pursuers fail to catch the evader in PP and CB methods, while our method can pass. PP was not able to catch the evader in half of the experiments and CB was not able to catch the evader in 75% of experiments. A key factor in our method is the rate of measurement coming from pursuers. In the experiments, it is set to 0.2 Hz for our proposed method, while in other methods it is 1 Hz. This means that the proposed method does not need to have measurements in all time steps and can lead to less memory consumption and computation overhead.
IV-A2 Trajectory
@@
To investigate the effect of different trajectories on the performance of the proposed method, various trajectories were tested. As shown in Fig. 5, the pursuers demonstrate accurate position estimation and effective pursuit of the evader. In the figure, the initial positions of the pursuers are indicated by circles, and there are four pursuers in the scene. Notably, the speed of the pursuers is twice that of the evader $(V_{p}=2V_{q})$.
Figure 5: The effect of different trajectories on the estimation over 500 time-steps has been investigated. The circle shows the start position of the robots. Pursuers try to get closer to the evader and catch it.
IV-A3 Number of Robots

### v2

@@ 3.3.1 Dynamic Factor of Evader
[…]
where, $||.||$ denotes the Mahalanobis distance [42], and $\Omega_{t}^{q}$ represents the information matrices of the variables. In Eq. (11), $x^{q}_{t}$ represents the evader’s position at time $t$. The cost function measures the prediction error of the relative transformation between two consecutive evader poses, comparing the predicted motion based on current pose estimates with the actual odometry measurements. This error is weighted by the uncertainty of those measurements.
3.3.2 Dynamic Factor of Pursuers
@@
3.3.3 Measurement factor between each pursuer and evader
@@
3.3.4 Measurement Factor between Pursuers and Obstacles
Pursuers need to localize themselves in the scene. To do this, the assumption is that a map of the environment is available for the pursuers in advance, and the pursuers utilize the map, along with the real-time measurement of the obstacles, to localize themselves. The static obstacles’ locations are assumed to be known and are used as landmarks for the localization of the robots. However, the target of the evader, as well as its position and orientation, are assumed to be unknown to the pursuers. The localization is performed via the proposed factor graph.
It is assumed that the pursuers have full observability of the scene and know the poses of the obstacles. They need to localize their own poses with respect to the obstacles. This measurement factor can be defined as follows:
[…]
where $\hat{x}^{p}_{t}$ represents the estimated position of a pursuer at time $t$, $x^{o}$ represents the position of obstacle $o$, and $h(\cdot)$ is the measurement model which measures the bearing and range between a pursuer and obstacles [11].
3.3.5 Planning Factor for Pursuers
@@
[…]
3.3.6 Collision Avoidance Factor for Pursuers
@@
where the objective is to define a safety bubble for each pursuer. The radius of the safety bubble is given by $d_{s}$. If the distance between pursuers $p_{i}$ and $p_{j}$ is less than $d_{s}$, the error term $1-\frac{d_{p}}{c_{1}}||$ is applied. Here, $c_{1}=2r_{R}+\epsilon$, where $\epsilon$ is a small safety distance and $r_{R}$ represents the robot’s external body radius (see Table 2 for values), and $d_{p}$ represents the distance between the two pursuers, $p_{i}$ and $p_{j}$.
3.3.7 Obstacle Avoidance Factor for Pursuers
@@
Here, $d_{o}$ represents the distance between each moving pursuer and the obstacles in the scene, while $c_{2}=r_{R}$ (see Table 2 for values). The optimized cost function generates the path based on the estimated position of pursuers ($\hat{x}^{p}_{t}$) and the evader ($\hat{x}^{q}_{t}$) at time $t$.
3.4 Time Complexity Analysis
[…]
The overall complexity is
@@ 3.4 Time Complexity Analysis
Figure 2: The number of different factors at each time step is shown in the figure above. In this figure, $f^{op}$ and $f^{mo}$ are increasing at the same rate, while $f^{dp}$, $f^{mp}$, and $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f^{movp}$} are increasing at the same speed.
As shown in Fig. 2, the number of measurement factors between pursuers and obstacles and obstacle avoidance factors increase faster compared with other factors and variables. The LM solver can be approximated by $\widetilde{O}(\epsilon^{-2})$ [46] where $\widetilde{O}$ exhibits the existence of logarithmic factors in $\epsilon$, and $\epsilon$ represents an $\epsilon$-stationary point. By considering $m$ as solver parameters and considering the post-optimization that takes place in $O(I)$, the overall complexity would be $\widetilde{O}(n\epsilon^{-2}(n^{2}+m)+nI)$.
3.5 Convergence Analysis
@@
$f_{i}(X_{i})\propto\exp\left\{-\frac{1}{2}\left\|h_{i}(X_{i})-z_{i}\right\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Omega}}^{2}\right\}.$ (19)
@@
$X^{\text{MAP}}=\arg\min_{X}\sum_{i}\left\|h_{i}(X_{i})-z_{i}\right\|_{\Omega_{i}}^{2}.$ (20)
The LM optimizer can solve Eq. (20) by linearizing it as:
$\Delta^{*}=\arg\min_{\Delta}(A\Delta-b)^{\top}(A\Delta-b).$ (21)
In Eq. (21), $A$ is the single measurement Jacobian, and $b$ shows all the prediction errors in the form of $z_{i}-h_{i}(X^{0}_{t})$. $\Delta$ denotes the state increment, the change in the evader’s pose at time $t$.
As a simple scenario, if there are two pursuers, $p_{1}$ and $p_{2}$, and they observe the evader only once the factor graph simplifies as shown in Fig. 3. This graph has a tree structure, and it converges surely [48]. In the factor graph, the factor function for the movements of pursuers is denoted as $f_{1}^{dp_{1}}$ for $p_{1}$ at time-step 0, as shown in Fig. 3. The prior factor of the pursuer $p$ is represented as $f_{0}^{p_{1}}$ at time-step 0. Variable elimination is an exact inference method. As an example, the probability function of variable $x_{3}^{q}$ can be found by summing out all other variables in the joint distribution:
@@
$\sum_{x_{0}^{p_{2}}}\cdots\sum_{x_{3}^{p_{2}}}P(x^{q}_{0},\ldots,x^{q}_{3},x_{0}^{p_{1}},\ldots,x_{3}^{p_{1}},x_{0}^{p_{2}},\ldots,x_{3}^{p_{2}}).$ (22)
In Eq. (22), all variables are marginalized out, except $x_{3}^{q}$. By starting the elimination process in the graph shown in Fig. 3, we begin by eliminating node $x_{0}^{p_{1}}$. The involved factors are $f_{0}^{p_{1}}$, $f_{0}^{mp_{1}}$}, and $f_{1}^{dp_{1}}$}. After eliminating $x_{0}^{p_{1}}$, two marginal factors are created:
@@
3.6 Uncertainty Minimization Analysis
@@
$x_{t}^{q}=x_{t-1}^{q}-(J^{T}J+\lambda I)^{-1}J^{T}e_{t}^{dq}.$ (30)
In the above equation, $J$ represents the Jacobian matrix of the cost function, and $\lambda$ is the damping factor that balances between the gradient descent and Gauss-Newton methods. By considering that the only factor existing in the problem is the dynamic factor, the uncertainty can be expressed as:
$\sigma_{T}=\sigma_{0}+\sum_{i=0}^{T}e_{t}^{dq}-\Delta_{LM}$ (31)
@@
4 Experiments
Experiments are conducted in two environments: (i) a simulated environment and (ii) a real-world hardware environment. The performance analysis is based on the average distance traveled by pursuers, time, and the number of time steps. In our study, to ensure fairness and consistency, we employed an exhaustive search over the parameters and weight space for the same scenario. This approach allows us to rigorously evaluate different weight configurations and verify that the reported results are not an artifact of arbitrary parameter selection, but rather reflect robust performance across a range of weight choices. In total, we evaluated 1000 combinations of input sets. In order to consider various scenarios, we vary the values related to dynamic factor uncertainties, measurement uncertainty, and the frequency of measurements.
To evaluate the performance of different weight settings, we measured the time required for four pursuers to capture an evader moving from left to right.
Figure 4: Time to capture the evader for four pursuers in 1000 iterations to find the optimal weights for our problem.
As shown in Fig. 4, the optimized weights catch the evader in 5 seconds, which is the minimum time to capture the evader over 1000 iterations. In order to show the correlation between each parameter and time to capture the evader the correlation analysis was conducted.
Figure 5: Correlation analysis between parameters and time to capture the evader (See Table 2 for parameters).
In Fig. 5, the correlation between the time to capture the evader and the parameters is shown. The correlation quantifies the strength and direction of the linear relationship between each parameter and the capture time. It is obtained by computing the Pearson correlation coefficient [49] across multiple simulation runs, where each run varies one or more parameters while recording the corresponding time required to capture the evader. A positive correlation indicates that increasing the parameter tends to increase the capture time, while a negative correlation indicates that increasing the parameter tends to reduce the capture time. The obtained optimal values utilizing exhaustive search are as follows:
  • 1.
    Dynamic factor: In this problem, the uncertainties for each step along the x-axis and y-axis are denoted as $\sigma_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}dx}}$ = 0.1 and $\sigma_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}dy}}$ = 0.1, respectively. Additionally, the uncertainty in orientation is $\sigma_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d\theta}}$ = 0.01.
  • 2.
@@ 4 Experiments
    Collision avoidance factor: In order to prevent pursuers from having collisions, this factor is utilized. The uncertainties for each step along the x-axis and y-axis are denoted as $\sigma_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}cpx}}$ = 1 and $\sigma_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}cpy}}$ = 0.1, respectively.
  • 5.
@@
    Obstacle avoidance factor: In order to prevent pursuers having collision with obstacles, this factor is utilized. The uncertainties for each step along the x-axis and y-axis are denoted as $\sigma_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}opx}}$ = 1 and $\sigma_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}opy}}$= 0.1 respectively.
Table 2 lists the tuned parameters used for the simulated experiments.
Variable Definition Value
$r_{R}$ Pursuer external body radius 0.3 m
$c_{1}$ $c_{1}=2r_{R}+\epsilon$ 0.61 m
$\epsilon$ Safety Margin 0.1 m
$c_{2}$ $c_{2}=r_{R}$ 0.3 m
$d_{s}$ Safety distance bubble 0.6 m
$\sigma_{dx}$ Uncertainty along x-axis in dynamic factor for pursuers and evader 0.1 m
$\sigma_{dy}$ Uncertainty along y-axis in dynamic factor for pursuers and evader 0.1 m
$\sigma_{d\theta}$ Orientation uncertainty in dynamic factor for pursuers and evader 0.01 rad
$\sigma_{\text{range}}$ Range measurement uncertainty in measurement factor 10 m
$\sigma_{\text{bearing}}$ Bearing measurement uncertainty in measurement factor 0.05 rad
$\sigma_{cpx}$ Uncertainty along x-axis in collision avoidance factor 1 m
$\sigma_{cpy}$ Uncertainty along y-axis in collision avoidance factor 0.1 m
$\sigma_{opx}$ Uncertainty along x-axis in obstacle avoidance factor 1 m
$\sigma_{opy}$ Uncertainty along y-axis in obstacle avoidance factor 0.1 m
Table 2: List of parameters used for the experiment.
4.1 Simulation
@@
The simulated environment is developed in a way that the number of pursuers, evaders, and obstacles in the scene can be modified. The simulated environment is developed in a $3500\times 3500$ window size. Additionally, the simulated environment is capable of simulating different trajectories of the evader. The simulation is implemented in C++ and OpenCV library is utilized for visualization purpose.
4.1.1 Comparison
@@
Various approaches have been explored in pursuit of strategies for capturing and chasing evaders. Among these, Qin et al. [13] introduced a hierarchical network-based model aimed at devising a cooperative pursuit strategy. It aims to solve the pursuit-evasion problem by generating utility trees (GUT) [13]. To ensure a fair comparison, the initial poses and speeds of both the evader and pursuers are kept the same for the proposed and compared methods. Two additional tactics for pursuers are also considered: Constant Bearing (CB) and Pure Pursuit (PP) strategies [12]. In the CB strategy, the bearing angle between a pursuer and the evader remains fixed until the time of capture. Conversely, in the PP strategy, the pursuer’s velocity vector aligns with the line of sight. Our proposed method operates at a frequency of 20 Hz on average, while the compared methods operate at a frequency of 30 Hz on average.
@@
Figure 6: Average distance traveled by pursuers (left) and time to capture the evader (right) for different approaches. $V_{p}$ and $V_{q}$ show the rate of pursuer speed in comparison to evader speed. PP and CB methods are not able to catch the evader at all speed rates, in which no values are reported for them.
[…]
4.1.2 Different Trajectories for Evader
@@
To investigate the effect of different trajectories on the performance of the proposed method, various evader trajectories were tested. As shown in Fig. 7, the pursuers demonstrate accurate position estimation and effective pursuit of the evader. In the figure, the initial positions of the pursuers are indicated by circles, and there are four pursuers in the scene. The measurement frequency is 1 Hz. The measurement frequency refers to the number of timesteps at which measurements are taken; a lower frequency means that measurements are obtained less often, which can increase uncertainty in the estimation process.
    Pursuer 1     Evader ground truth
    Pursuer 2     Evader estimate
    Pursuer 3     Evader start point
    Pursuer 4     Evader final point
    Pursuer start point     Obstacles
Figure 7: The effect of different trajectories on the estimation over 1600 time-steps has been investigated. The black circle shows the start position of the pursuers. They try to get closer to the evader and catch it. Dashed lines show their trajectory. The dotted black line represents the predicted trajectory of the evader, and the red line represents the evader’s actual (ground truth) trajectory. Obstacles are shown as gray squares, and the cross marks the evader’s final position. Utilizing FG-PE, the evader is captured when operating at a measurement frequency of 1 Hz.
To demonstrate that the proposed method is able to solve the problem even when the measurement frequency is low, the method was evaluated at a measurement frequency of 0.2 Hz.
Pursuer 1 Evader ground truth
Pursuer 2 Evader estimate
Pursuer 3 Evader start point
Pursuer 4 Evader final point
Pursuer start point Obstacles
Figure 8: Performance when the frequency of measurement is 0.2 Hz.
This reduces computation time, as the frequency of measurement is decreased. As shown in Fig. 9, when the measurement frequency between pursuers and the evader decreases, the average time per iteration also drops.
Figure 9: Average time of iteration for different frequency of measurement between pursuers and evader.
Another experiment was conducted to show that, even when pursuers have different levels of odometry noise, they can still estimate the pose of the evader and plan to catch it.
Pursuer 1 Evader ground truth
Pursuer 2 Evader estimate
Pursuer 3 Evader start point
Pursuer 4 Evader final point
Pursuer start point Obstacles
Figure 10: Performance when pursuers have different levels of odometry sensor noise. The shaded circle around each pursuer’s start point indicates the level of noise, representing the noise covariance matrix.
As shown in Fig. 10, although the pursuers have different levels of odometry sensor noise, the proposed approach allows them to estimate the evader’s pose and plan to catch it.
4.1.3 Number of Robots